A body rolled on a surface with a speed of 10m/sec comes to rest after covering a distance of 50m.Find the coefficent of friction.g=10m/sec2

Grade:11

1 Answers

Arun
6 years ago
Dear nandini
 
v^2 - u^2 = 2as 

0-100=2*a*50 

-100/100=a 

a=-1 m/s^2 
Negative sign shows retardation. 

let n be the coeff. of friction, 

n*mg=m*a 
n=a/g 
n=1/10

n=0.1
 
Regards
Arun(askIITians forum expert)